OpenStudy (anonymous):

CAN ANYBODY PLEASE HELP ME ALREADY? Celine has a bottle that contains 20% of milk and the rest water. The bottle has 1 liter of water. Write an equation in one variable that can be used to find the total number of liters of milk and water in the bottle. Define the variable used in the equation.

OpenStudy (anonymous):

I would like to be walked through it mostly, not just told the answer.

OpenStudy (anonymous):

The problem states that only one variable can be used. Let L be the total number of liters of water and milk combined. 20% of the total is Milk. Therefore, Milk = L x 20/100 = 0.2 L The rest of it is Water. So Water = L - 0.2L = 0.8L The problem states the bottle has 1 liter of water. So 0.8L = 1 Answer to a) 0.8L = 1 is the equation where L is the total number of liters of water and milk combined. Milk = 0.2L and Water = 0.8L b) Solve for L in 0.8L = 1 L = 1/0.8 = 1.25 Milk = 0.2L = (0.2)(1.25) = 0.25 liters

OpenStudy (anonymous):

@Prestianne nice job

OpenStudy (anonymous):

:D thanks!

OpenStudy (anonymous):

Yes, thank you.

OpenStudy (anonymous):

:) ur welcom

OpenStudy (anonymous):

See, I found the answer without having to make an equation. But how to actually write the equation itself is what was confusing me.

OpenStudy (anonymous):

ohh, well now i hope u understand :)

OpenStudy (anonymous):

Actually, it didn't help me a single bit. I thought it did.... but I still don't get it.

OpenStudy (anonymous):

It'd not that what you said is wrong or anything, I am just having trouble comprehending it.